Kindly answer these problems please because i am still quite confused in answering these problem using system of linear equations.
2) A diving pool can be filled by a pipe in 10 hrs. if the drain is open. If the pipe was allowed to run with the drain for 2 hrs. and the drain was closed, it would require the pipe 4 more hrs. to fill the pool. How long would it take the pipe to fill the pool if the drain is closed from the start?

Thank you in advance! Hoping for your immediate reply.

Answer by sudhanshu_kmr(1152)


let assume that pipe fill the tank in x hours,so in one hour 1/x part of tank.
and drain empty the tank in y hours, so in one hour 1/y part of tank.
here obviously value of x is less then y.
portion of tank filled in one hour when both are open = 1/x - 1/y
according to question,

10(1/x -1/y) = 1 => 10y - 10x = xy
and 2(1/x -1/y) + 4* 1/x = 1
from first equation we find, y = 10x / (10-x)
now, put this value of second equation,
2( 1/x - (10-x)/10x )+ 4*1/x = 1
=> 2 * 1/10 + 4/x = 1
after solving, x= 5, y = 10
x = 5 hours
so, if drain is closed the pipe will fill the pool in 5 hours...


if any problem, welcome to contact.....